A loan of 10000 is to be repaid over 10 years by a level annuity payable monthly in arrears. The amount of the monthly payment is calculated on a basis od an interest rate of 1% per month effective,. For which monthly repayment the capital first exceeds the interest


1
Expert's answer
2021-12-12T17:01:01-0500

"10000=R\frac{1-1.01^{-10\cdot12}}{0.01}"


monthly repayment:

"R=143.47"


"FV=\frac{R((1+i)^n-1)}{i}=10000"


monthly repayment when the capital first exceeds the interest:


"n=ln(\frac{10000i}{R}+1)/ln(1+i)=53.15" months


Answer: 54-th month